OpenStudy (anonymous):

check my answers What are the solutions of the quadratic equation? x2 + 11x = —24 (1 point) *–3, –8 –3, 8 3, –8 3, 8

OpenStudy (anonymous):

2. What are the solutions of the quadratic equation? 4x2 + 34x + 60 = 0 (1 point) *6, 2/3 -5/2,-2/3 –6, –5/2 –6, 4 3. Solve by using tables. Give each answer to at most two decimal places. –2x2 – 4 = –8x (1 point) 0.59, 3.41 –3.41, –0.59 *1.17, 6.83 –1.41, 1.41

OpenStudy (anonymous):

@AriPotta @NickR @tkhunny

OpenStudy (tkhunny):

What's your plan? Please show your work.

OpenStudy (anonymous):

im not sure i pretty much just try to go through them and geussed bcuz i wasnt sure im 90% sure i got them all wrong

OpenStudy (tkhunny):

That's horrible. We'll need better than a 10% confidence ratio, right? x^2 + 11x = —24 x^2 + 11x + 24 = 0 (z+8)(x+3) = 0 Now what? 4x^2 + 34x + 60 = 0 2x^2 + 17x + 30 = 0 (2x + _____)(x + _____) = 0 Now what? –2x^2 – 4 = –8x –2x^2 + 8x – 4 = 0 x^2 - 4x + 2 = 0 Quadratic Formula? Be neat. Be consistent. Be thorough. No guessing.

OpenStudy (anonymous):